Recent Announcements


LUCI-AO Offline Until Further Notice

posted Jul 5, 2021, 9:51 AM by Barry Rothberg

Operations with the LUCI-1 and the SX Adaptive Secondary have been suspended for the remainder of 2021A.  This is due to a safety issue with the Adaptive Secondary.  Seeing limited observations are not affected.  The observatory is working to address the issue with the goal of having LUCI-1 AO available again by the start of 2021B.

LUCI AO Webpages Revamp Begins

posted May 13, 2021, 1:21 PM by Barry Rothberg   [ updated May 13, 2021, 1:24 PM ]

As part of the upgrades/improvements to AO at LBTO, the LUCI AO webpages are being reorganized.  A new Phase 1 Planning Page is available for diffraction limited AO.  These pages are accessible from the sitemap on the righthand sidebar.

MODS2 Dichroic Mechanism [repaired 11-March-2021]

posted Mar 5, 2021, 6:32 AM by Olga Kuhn   [ updated Mar 11, 2021, 4:19 PM ]

At the beginning of March, the MODS2 dichroic was failing and it was found that the gear was stripped. A replacement part is on the way and we hope that the repair will be made by Wednesday, March 11.

Meanwhile, the dichroic drum has been fixed at the dichroic to allow dual channel configurations: Dual Grating, Dual Imaging and Dual Prism; and the script engines have been made to override the scripts when they call for a different configuration in the "instconfig" (in the case of observation scripts) or ACQCamera (in the case of acquisition scripts).

PIs do not need to change their acquisition scripts, and they need to be aware that observations will be limited to dual mode.

A two-day repair finished today, Thursday, 11-March (in the above note, the 11 was correct but the day of the week was not --- at the time of writing, I had thought it was a one-day repair and would be done by the start of the 11-March partner block). Thanks for Elliott, John and the MODS team at OSU for getting it back into operation.

LUCI-2 Filter Wheel 1 Issue

posted Jan 20, 2021, 1:54 PM by Barry Rothberg   [ updated Jan 20, 2021, 1:55 PM ]

We strongly recommend PIs to check scripts to avoid using HeI+J for Telluric crossed filters. Please use OH1060+z for Telluric crossed filters.  At least one of the switches in the filter wheel which holds the narrow and medium band filters (as well as clear and  blind) is showing a problem.  The filters potentionally affected are:  P-Beta, J-low, J-high, PVlens (not used for science), OH 1190, FeII, H2, HeI.

LUCI-1 Bad Amplifier Channel Recovered

posted Sep 6, 2020, 5:12 PM by Barry Rothberg   [ updated Sep 6, 2020, 5:13 PM ]

After summer shutdown reparirs, the bad amplifier channel on the LUCI-1 detector is functioning now.


MOS Mask Deadline Extended to August 15, 2020 for 2020B

posted Jul 24, 2020, 2:07 PM by Barry Rothberg   [ updated Jul 24, 2020, 2:08 PM ]

Due to the current set of circumstances, we are extending the MOS mask deadline to August 15, 2020 for the 2020B semester.  Masks submitted in time for the original deadline are appreciated. Please contact slitmasks@lbto.org for any questions regarding mask design or deadlines.

New Version of MMS is Released!

posted Jul 24, 2020, 1:57 PM by Barry Rothberg   [ updated Jul 24, 2020, 1:58 PM ]

Version 2.0.1 of the MMS software is now released to the LBT community.
Information on installation and use can be found here:
https://sites.google.com/a/lbto.org/mods/preparing-to-observe/mask-preparation-2
Starting with 2021A, ALL new MODS MOS masks must use the new version of MMS.  We are accepting MODS MOS masks designed with V2.0 and V1.60 for 2020B, but users are STRONGLY encouraged to use the new version. 

Updated MODS1+SX & MODS2+DX zero points

posted Apr 24, 2020, 12:58 PM by Olga Kuhn   [ updated Apr 24, 2020, 1:02 PM ]

The MODS Sensitivity page has been updated to give current zeropoints for direct and dual grating modes of MODS1 and MODS2. Since MODS1 was commissioned, the sensitivity of the telescope+instrument combination has declined and so updated zeropoints were in order. The updated zeropoints for MODS1+SX and MODS2+DX are based on measurements made in September 2019 and are consistent with data taken up until the spring of 2020. SNR tracks which use these updated zeropoints have been posted for observations at g (blue channel) and r, i and z (red channel).

The first update was made on April 7, but an error was found in the MODS1B zeropoints that had been entered into the table (but not the ones used for the SNR tracks), and the table was corrected on April 23.

Update to LUCI Headers - DIMM Values

posted Mar 5, 2020, 12:23 PM by Barry Rothberg   [ updated Mar 5, 2020, 12:33 PM ]

The DIMM header value for LUCI-1 and LUCI-2 has been updated. As of UT 05-March-2020 the header now shows:
DIMMFWHM=                  -1. / Zenith-corr FWHM(arcsec)@0.5micron
(on-sky observations will show values other than -1).  The values going forward reflect the astronomical standard of the DIMM FWHM (in arcseconds) value corrected to zenith and a wavelength of 0.5 microns.  This allows for the DIMM FWHM to then be computed to the elevation (airmass) of the astronomical target at the wavelength of the observations.

Prior to this update, the DIMM value in the headers:
DIMMFWHM=                  -1. / ARGOS DIMM FWHM seeing
reflected the FWHM measured at the elevation of the DIMM camera, which may differ significantly from the elevation of the science target.  Without the elevation of DIMM camera, this value cannot be corrected to the elevation and wavelength of the science observations.  Users should *not* use the DIMM FWHM values in the headers prior to UT 05-March-2020 as a quantitative measurement of the seeing of the science observations.    However, any DIMM FWHM data recorded from the FACSUM during observations (or with LBTplot) do reflect the astronomical standard value corrected to zenith and a wavelength of 0.5 microns.

LUCI-1 Update & Detector Issue

posted Feb 23, 2020, 7:15 PM by Barry Rothberg   [ updated Mar 12, 2020, 6:25 PM ]

LUCI-1 Detector Issue:  After a successful repair of the detector focus block a new issue was found with the detector. One of the 32 channels on the detector is reading out only 0 values.  LUCI-1 is available for Seeing Limited and ESM imaging and spectroscopcy - but users should take into account the impact of the loss of this channel.  The images below provide information on the angular size and location of the issue.  The fits file available below shows the readout issue and can be used to help plan changes to LUCI-1 programs.
LUCI-1 amplifier channel issue


Longslit:
The image below compares the dead channel on a dispersed image with the N1.8 camera with a longslit dither using a +/-50" dither.  This indicates that a + dither of ~40-60" (plus a few arcseconds depending on the accuracy of centering the target in the detector in DET Y and any movement in the Fold Mirror #4, and some difference between where the G210 and G200 grating disperses light) will land a longslit spectrum in the dead amplifier channel. 


The two files below represent an N3.75 acquisition image (luci1.20200306.0172.fits) and a G200 Neon arc lamp using a 1" slit and the N1.8 camera (luci1.20200306.0192.fits).  These can be used to estimate the impacts to science imaging/acquisition as well as spectroscopy.  Additional files will be made available in other modes (e.g. N1.8 G210 spectra, N3.75 G210 spectra).


MOS Mask:
The example below shows how the change in camera between acquisition (N3.75) and grating with the N1.8 camera is affected by the bad amplifier channel.  The left two panels show the dispersed spectrum of the MOS mask (top) and what the MOS mask looks like with the N1.8 camera.  The right image show what the MOS mask looks like during an acquisition with the N3.75 camera. 
  • On the right, the bad channel affects the slit below the wide slit, ~87.5" below the top of the detector (the line marks the distance between an alignment hole and the bad channel). 
  • On the left, the change in camera and grating changes things.  The bad amplifier channel now overlaps with the wide slit and is ~58.7" below an alignment hole.
  • PIs must consider the impact of the bad channel on the acquisition and alignment and then how this shifts if the program uses the N1.8 camera for spectroscopcy (the difference in where the G200 and G210 disperse light on the detector is shifted slightly ~15-20 pixels).
  • If the spectroscopy uses the N3.75 camera, then where the bad channel affects the mask in the acquisition image is approximately where it affects the dispersed light.
  • PIs can use the LMS software to estimate the impact during acquisition and the impact to spectroscopic observations. 
    • Load the image used to make the mask (or a DSS image corresponding to the coordinates of the field.
    • Load the LMS file.
    • Use the right button on the mouse to draw a vector.  The numbers of the vector represent either arcseconds (up to 59.99) or arcminutes.  Pick one of the alignment boxes along the top of the mask (there are 6) and draw a vector with lengths equal to 1' 26.5"  and then 1' 35.1"  - this corresponds to the width of the bad amplifier channel on the N3.75 camera.  Note if any alignment stars are affected.
    • If using the N1.8 camera, draw a vector from one of the alignment boxes of lengths equal to 58.7" and then ~1'14".  This corresponds to the width of the bad amplifier channel on the N1.8 camera.  Note if any of the science slits fall within this range.  Please allow a little wiggle room depending on FM4 mirror alignment and slight shifts in how light is dispersed between the G200 and G210 gratings.
    • Also keep in mind that there is a slight tilt in the dispersed light from each grating (can be seen in the MOS and Longslit examples here).  This tilt means that the bad amplifier may cut through only part of the slit, or also affect a small portion of a neighboring slit's dispersed light. To err on the side of caution one may assume ~20" high zone of avoidance on the N1.8 camera and ~10" zone of avoidance for spectroscopcy with the N3.75 camera.

Comparions MOS mask N1.8 and N3.75



1-10 of 43